OpenStudy (anonymous):

Help please. Which set of numbers does not create a triangle? 900 cm, 60 cm, 0.5 m 12 m, 34 m, 40 m 145 ft, 14 yd, 105 ft 16 yd, 12 yd, 27 yd

OpenStudy (anonymous):

I was thinking that the correct answer is the 2nd one but not sure

OpenStudy (tkhunny):

From each set of three length, ANY two must sum to more than the third. All three sets must work. Example 3-4-5 is a triangle 3+4> 5 3+5> 4 4+5> 3

OpenStudy (anonymous):

ok

OpenStudy (tkhunny):

Let's try #2 12 m, 34 m, 40 m 12 + 34 = 46 > 40 12 + 40 = 52 > 34 34 + 40 = 74 > 12 Looks like a triangle, to me.

OpenStudy (anonymous):

ok I see how you did that

OpenStudy (tkhunny):

Careful on #1 and #3. The units are not consistent. Note: It takes THREE successes to make a triangle. It takes only one failure to rule it out.

OpenStudy (anonymous):

Ok yes I was looking at the units that are not the same

OpenStudy (tkhunny):

#3 145 ft, 14 yd, 105 ft Forget that and work with this: 145 ft, 42 ft, 105 ft

OpenStudy (anonymous):

145 + 42 = 187 > 105 145 + 105 = 250 > 42 42 + 105 = 147 > 145

OpenStudy (anonymous):

So it would make a triangle correct?

OpenStudy (tkhunny):

Okay. Two more to try.

OpenStudy (anonymous):

Ok so I just did the last one it makes a triangle too, so the answer it #1 that does NOT make a trirangle.

OpenStudy (tkhunny):

Prove it! 60 + 500 = 560 < 900. Done!!

OpenStudy (anonymous):

I just worked it out too, thank you so much for explaining this to me and helping me.

OpenStudy (tkhunny):

Good work!